AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 108     Entry time: Tue Oct 5 17:30:01 2021
Author: CB RSS 
Subject: NEG activation tests 
Found D2000 #1 activated and at 250C (25W) 
to prevent getting dirty when other NEGs 
are activated. See previous entries.

Attempted activation of D2000 #2 using T 
390C , rise time 30 m, hold time 90 m

After 30 m temperature inside chamber 
increased by 5C. 390 C corresponds to 75 
W.

Attempted activation of D2000 #3 while #2 
is holding temperature. Same parameters as 
above except T 400.

After 30 min, temperature in chamber up by 
10 more C to 110.

Attempted activation of #4 as above.

Set #2 then #3 then #4 to 250 C after hold 
time is over to protect them. All D2000 
activated. Max temperature on Kapton 
bundle 110 C.



Attempted activation of 2xUHV1400 linked 
to top C4, channel 2.

No thermocouple available. Used power 200 
W (2x100) rise time 30 m and hold time 90 
min. Upon reaching 200 W no effect on 
chamber interior temperature. Increased to 
300 W. Temperatures increased to 120C max 
roughly. Acceptable.

Had to decrease hold time to 60 m to avoid 
remaining in cave alone for too long. 
Likely okay with 300 W (450 C?). Left on 
150W (400C?) to protect from subsequent 
activations tomorrow.

Because cables are unlabeled it's not 
clear how far the activated UHV1400 are 
from the detectors. If close, this recipe 
works. If far, we may have to use lower 
power for other modules.
Attachment 1: 20211005_131259.jpg  3.253 MB  Uploaded Tue Oct 5 18:45:38 2021  | Hide | Hide all
20211005_131259.jpg
Attachment 2: 20211005_171403.jpg  2.769 MB  Uploaded Tue Oct 5 18:46:23 2021  | Hide | Hide all
20211005_171403.jpg
Attachment 3: 20211005_161317.jpg  2.735 MB  Uploaded Tue Oct 5 18:46:54 2021  | Hide | Hide all
20211005_161317.jpg
ELOG V3.1.4-unknown